Given$100,000to​ invest, construct a​ value-weighted portfolio of the four shares listed below

Stock   Price/Share ($)   Number of shares outstanding (millions)
Golden Seas   $11   1.04
Jacobs and Jacobs   $24   1.57
MAG   $47   31.77
PDJB   $11   13.41

Enter the portfolio weights below.  ​(Round to three decimal​ places.)

Share

​% of total value

​(portfolio weight)

Golden Seas

nothing​%

Enter the portfolio weights below.  ​(Round to three decimal​ places.)

Share

​% of total value

​(portfolio weight)

Jacobs and Jacobs

nothing​%

Enter the portfolio weights below.  ​(Round to three decimal​ places.)

Share

​% of total value

​(portfolio weight)

MAG

nothing​%

Enter the portfolio weights below. ​(Round to three decimal​ places.)

Share

​% of total value

​(portfolio weight)

PDJB

nothing​%

Enter your answer in each of the answer boxes.
Answer #1
Company Market value market value weights
Golden seas 11*1040000=11440000 11440000/1689820000= .677
Jacobs and Jacobs 24*1570000=37680000 37680000/1689820000= 2.230
MAG 47*31770000=1493190000 1493190000/1689820000= 88.364%
PDJB 13410000*11=147510000 147510000/1689820000= 8.729%
Total 1689820000 100%